OpenStudy (anonymous):

Find the Orthogonal Projection of f onto g. Use the inner product in C[a,b] C[-pi, pi] f(x)=x and g(x)= cos 2x

OpenStudy (anonymous):

if i remember correctly this is \[\int_{-\pi}^{\pi}x\cos(2x)dx\]

OpenStudy (anonymous):

yes that is correct but i am having difficulties integrating

OpenStudy (anonymous):

oh parts for this one. pretty sure it is zero

OpenStudy (anonymous):

oh wait it is zero from your eyeballs because \[h(x)=x\cos(x)\] is odd and you are integrating from \[-\pi\] to \[\pi\]

OpenStudy (anonymous):

 This is the formula